find area of triangle
1 2/5 yrds by 4 1/2 yrd

Answer by hyperasad(6)

area of triangle = 1/2 x height x base
in this case
height = 1 2/5 = 7/5
base = 4 1/2 = 9/2
placing these values in formula
AREA = 1/2 x 7/5 x 9/2
= 63/20
= 3 3/20 sq. yards